JET SPECTROSCOPY OF THE H-F OUT-OF-PLANE LIBRATIONAL FUNDAMENTAL BAND OF HY-DROGEN FLUORIDE PENTAMER IN THE 741 cm^{−1}REGION

\underline{THOMAS A. BLAKE}, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, 3020 Q Avenue, Richland, WA 99352 (PNNL is operated for the US Department of Energy by the Battelle Memorial Institute under contract DE-AC06-76RLO 1830).

Using our pulsed slit-jet, infrared diode laser spectrometer, we have observed a rotationally-resolved parallel band centered at 741 cm^{−1}in a molecular beam formed by expanding a mixture of 4% HF in helium through a 200 μm x 12 cm slit with a backing pressure of ca. 1000 Torr. A set of excited and ground state spectroscopic constants were fit simultaneously to a data set of 75 transitions using a symmetric top, rigid rotor Hamiltonian. The fit constants are ν_{0} = 740.9696(7) cm^{−1}, B′ = 0.07507(5) cm^{−1}, (C′ − B′) − (C′′ − B′′) = 0.0014(1) cm^{−1}, and B′′ = 0.07570(5) cm^{−1}. Recently, we have reported the observation of two librational bands of HF tetramer in the 710 to 775 cm^{−1}region: one band is a perpendicular band centered at 752 cm^{−1}and the other band is a parallel band centered at 714 cm^{−1}. The ground state B value of the tetramer was found to be 0.132081(1) cm^{−1}. For hydrogen fluoride tetramer and pentamer, theory predicts (Maerker, et al.) ^{a} planar, oblate symmetric top structures at their global minima with B_{e} values of 0.136 and 0.0767 cm^{−1}for (HF)_{4} (C_{4h} symmetry) and for (HF)_{5} (C_{5h} symmetry), respectively. We have assigned the 741 cm^{−1}band as the (A′′) HF out-of-plane librational fundamental of the pentamer, the 714 cm^{−1}band as the (A_{u}) HF out-of-plane librational fundamental of the tetramer, and the 752 cm^{−1}band as the (E_{u}) HF in-plane librational fundamental of the tetramer.

aC. Maerker, P. R. Schlyer, K. R. Liedl, T. -K. Ha, M. Quack, M. A. Suhm, J. Comp. Chem. 18 1695 (1997).